### v1

# DAC: Decomposed Automation Correction for Text-to-SQL

Abstract
Text-to-SQL is an important task that helps people obtain information from databases by automatically generating SQL queries. Considering the brilliant performance, approaches based on Large Language Models (LLMs) become the mainstream for text-to-SQL. Among these approaches, automated correction is an effective approach that further enhances performance by correcting the mistakes in the generated results. The existing correction methods require LLMs to directly correct with generated SQL, while previous research shows that LLMs do not know how to detect mistakes, leading to poor performance. Therefore, in this paper, we propose to employ the decomposed correction to enhance text-to-SQL performance. We first demonstrate that decomposed correction outperforms direct correction since detecting and fixing mistakes with the results of the decomposed sub-tasks is easier than with SQL. Based on this analysis, we introduce Decomposed Automation Correction (DAC), which corrects SQL by decomposing text-to-SQL into entity linking and skeleton parsing. DAC first generates the entity and skeleton corresponding to the question and then compares the differences between the initial SQL and the generated entities and skeleton as feedback for correction. Experimental results show that our method improves performance by $3.7\%$ on average of Spider, Bird, and KaggleDBQA compared with the baseline method, demonstrating the effectiveness of DAC  [^1] Our code and data are released in https://github.com/zirui-HIT/DAC..
1 Introduction
Text-to-SQL is an important task, significantly reducing the overhead of obtaining information from the database by generating corresponding SQL based on user questions (Deng, Chen, and Zhang 2022). Specifically, about the text-to-SQL task, users provide the question and the related database, and then the model generates the corresponding SQL results. Currently, the methods based on Large Language Models (LLMs) have become the mainstream for text-to-SQL because LLMs can achieve higher performance than fine-tuned models, without fine-tuning (Hong et al. 2024b; Li et al. 2024a; Shi, Tang, and Yang 2024; Kanburoğlu and Tek 2024). Therefore, in this paper, we mainly focus on how to enhance the performance of the text-to-SQL task based on LLMs.
Recent research shows that automated correction is an effective method to improve text-to-SQL performance (Chen et al. 2024; Askari, Poelitz, and Tang 2024; Xie et al. 2024; Wang et al. 2024a), where LLMs first generate an initial SQL and then detect mistakes of the SQL as correction feedback (Pan et al. 2024). For instance, SQL-CRAFT (Xia et al. 2024) assesses whether the SQL execution results on databases satisfy the user question and corrects the SQL accordingly. EPI-SQL (Liu and Tan 2024) builds an example library of error cases and then retrieves similar cases from the library to guide each correction. However, previous studies indicate that it could be not effective for LLMs to directly correct all mistakes in the generated SQL since it is hard for LLMs to detect mistakes in the generated results (Zhang et al. 2024a; Kamoi et al. 2024; Zhang et al. 2024c).
Figure 1: The comparison between direct correction and decomposed correction. Direct correction shows poor performance since LLMs do not know how to detect mistakes. Decomposed correction brings better performance since it is easier to detect mistakes from the decomposed tasks of entity linking and skeleton parsing.
Previous works (Dhuliawala et al. 2024; Vacareanu et al. 2024) show that correction with the decomposed reasoning process has better performance than directly correcting because detecting and correcting the mistakes in the reasoning process is easier than in the final answer. Therefore, in this paper, we discuss that: (i) Experimentally, we present that correction with decomposed sub-tasks is more effective than direct correction; (ii) Methodologically, we propose correcting the text-to-SQL results based on the decomposed tasks of entity linking and skeleton parsing to enhance the performance of the automation correction.
First, we discuss that correction with the decomposed sub-tasks can bring better correction performance than the direct correction, as shown in Figure 1. We first present that decomposing into sub-tasks ensures better automation correction performance since mistakes can be more easily detected and corrected based on sub-tasks than using the original generation result. Applying the above conclusion, we propose decomposing text-to-SQL into two sub-tasks: entity linking and skeleton parsing, which are the two most crucial capabilities for addressing the text-to-SQL task (Li et al. 2023a; Deng, Chen, and Zhang 2022; Hong et al. 2024b).
Based on the above analysis, we propose Decomposed Automation Correction (DAC), which corrects the generated SQL with the mistakes of entity linking and skeleton parsing as feedback. Schema linking denotes detecting the table and column names related to the question, where we use LLMs to directly generate relevant entities based on the user question and the database, following Li et al. (2023c). Skeleton parsing denotes generating the SQL skeleton corresponding to the user question, where we employ LLMs to generate the SQL skeleton according to the prior research (Guo et al. 2024b, a). After generating the entities and skeleton, we take their inconsistencies with the initial SQL as feedback for the correction, as shown in Figure 1.
To validate DAC, we conduct experiments on three mainstream text-to-SQL datasets: Spider (Yu et al. 2018), Bird (Li et al. 2023b), and KaggleDBQA (Lee, Polozov, and Richardson 2021). The experiments show that our method achieves an average improvement of $3.7\%$ compared to baselines, and a $1.3\%$ improvement over previous text-to-SQL automation correction methods, demonstrating the effectiveness of DAC.
Our contributions are as follows:
  • We experimentally discuss that decomposed correction has better performance than direct correction;
  • We present DAC, which improves text-to-SQL automated correction performance based on the decomposed sub-tasks of entity linking and skeleton parsing;
  • Experiments show that DAC improves performance by $3.7\%$ compared to baselines and $1.3\%$ compared to the previous text-to-SQL automated correction methods, proving the effectiveness of our method.
2 Discussion
Figure 2: The illustration of our discussion, taking the question “Order the stock idx with earnings more than 5,000” as an example. About the direct correction (left part), it is challenging for LLMs to pinpoint specific mistakes. After decomposing the SQL into sub-tasks (right part), it is easier for LLMs to identify and correct the mistakes.
In this section, we discuss that decomposing text-to-SQL into sub-tasks can effectively enhance SQL correction performance. First, we demonstrate that correction by decomposing into sub-tasks is more effective than direct correction since it is easier for LLMs to detect and correct mistakes based on the sub-tasks. Then, we propose to correct the text-to-SQL results by decomposing the task into two sub-tasks: entity linking and skeleton parsing, thereby enhancing the SQL correction performance.
2.1 Performance of Sub-Task Correction
We consider the correction performance of task decomposition in two parts: correction results satisfy the sub-tasks, and correction results that satisfy the sub-tasks also satisfy the original task. We propose that decomposing the task into sub-tasks can make the correction results better satisfy the sub-tasks, while the correction results satisfying the sub-tasks could not satisfy the original task. The illustration of the above proposition is shown in Figure 2.
Intuitively, the original task can be viewed as a combination of multiple sub-tasks. Since the sub-tasks used are included in the original task, according to Wang et al. (2024b), generating correct answers to sub-tasks is easier than the original task, thereby it is easier to detect the mistakes based on these sub-tasks than the original task. However, because the used sub-tasks do not entirely encompass the original task, even if the correction results satisfy all the used sub-tasks, it cannot guarantee that the results fully meet the original task. Therefore, when decomposing sub-tasks, it is necessary to ensure that the sub-tasks can maximally cover the original task so that the correction results satisfying the sub-tasks are more likely to satisfy the original task.
2.2 Adaption to Text-to-SQL
We apply the discussion above to the text-to-SQL task, where we use entity linking and skeleton parsing as sub-tasks for automated correction. Entity linking refers to identifying relevant table and column names from the given database schema based on the user question (Liu et al. 2021). Skeleton parsing denotes creating the corresponding SQL skeleton of the user question, which removes the table names, column names, and values in SQL (Guo et al. 2024b). We use these two sub-tasks because previous work regards them as the most crucial capabilities for text-to-SQL (Li et al. 2023a; Deng, Chen, and Zhang 2022; Hong et al. 2024b).
While based on the discussion above, correction of the sub-tasks could not ensure the result correctness of the text-to-SQL task. Therefore, we statistic the proportion of the correct SQL in which the entity and skeleton are correct. The statistical results, as shown in Table 4, are very close to one hundred percent This result shows that the correction results in line with our proposed sub-tasks are also basically in line with the text-to-SQL task, demonstrating the effectiveness of our sub-task decomposition method for the text-to-SQL automated correction.
3 Methodology
Figure 3: The pipeline of DAC, which consists of five steps: (i) SQL Generation: Generate the initial SQL; (ii) Entity Linking: Detect the question-related entities; (iii) Skeleton Parsing: Generate the SQL skeleton of the question; (iv) Comparison: Determine the inconsistencies between the initial SQL and the linked entities and parsed skeletons as feedback; (v) Correction: Correct the SQL with the comparison feedback.
Following the discussion above, in this section, we present DAC, which generates automated correction feedback by decomposing the text-to-SQL task into the sub-tasks of entity linking and skeleton parsing. The illustration of DAC is shown in Figure 3, and the prompts of DAC are present in the supplementary material.
3.1 SQL Generation
Before the correction, we first generate an initial SQL to be corrected. Following previous work (Xia et al. 2024), we use few-shot learning to generate SQL. We use BM25 to select demonstrations similar to the user question from the given demonstration pool, together with the database and the user question as the input. With the input, LLMs directly generate a single SQL as the initial result to be corrected.
3.2 Entity Linking
After obtaining the initial SQL, we aim to identify mistakes in the linked entities. To achieve this, it is essential to know the entities related to the user question, specifically the table and column names in the question-related database. Following previous work (Andrew et al. 2024; Goel et al. 2023), we use LLMs to select the question-related tables and columns. We employ the tokenized user question and the database as inputs and output the entity linking information in the format of List[Dict[str, str]] in Python, where the length of the list matches the number of tokens in the tokenized question. We use this format because the performance of entity linking generated in programmatic format by LLMs is superior to only generating entities (Li et al. 2023c).
Following Liu et al. (2021), each Dict includes three keys: (i) token: the corresponding token in the original question; (ii) schema: the corresponding table name or column name in the database; (iii) type: including “tab”, “col”, and “val”, representing the table name, column name, or a condition value that the token corresponds to respectively.
3.3 Skeleton Parsing
After identifying the linked entities, we seek to parse the skeleton corresponding to the user question to identify mistakes in the initial SQL generation on the skeleton. Previous research indicates that LLMs can parse the SQL skeleton based solely on the user question since the skeleton corresponds to the syntactic logic of the question without requiring database information (Gu et al. 2023; Pourreza et al. 2024; Kothyari et al. 2023). Therefore, we use only the user question as the input and ask LLMs to parse an SQL skeleton corresponding to the question. Subsequently, we remove entities from the parsed SQL, since these entities could not correspond to the column names or table names in the original database, thus obtaining the SQL skeleton.
3.4 Comparison
After obtaining the linked entities and parsed skeleton, we directly compare and identify the inconsistency parts between them and the initial SQL without LLMs as the correction feedback. We first extract the entities and skeletons in the initial SQL according to the SQL syntax. For linked entities, we determine which linked entities are mentioned in the question but not used in the initial SQL. We do not consider entities that are present in the SQL but not mentioned in the question as mistakes because the user question could not explicitly mention some entities used in the SQL. For example, some tables are used to bridge the other two tables by the foreign keys without being mentioned by the user question. About the parsed skeleton, we correct the initial SQLs whose skeletons are inconsistent with the parsed skeletons. We input the completed parse skeleton as the feedback, where do not indicate specific mistaken keywords because different keywords could appear multiple times in the SQL, while too specific information could confuse the model about which keywords exactly need to be modified.
3.5 Correction
Upon constructing correction feedback, we use them to correct the initial SQL. If an error message occurs during SQL execution, we use the error message as feedback, following Chen et al. (2024). We take the database, user question, initial SQL, and feedback as input, and ask LLMs to output the corrected SQL. During the correction, we first correct entity mistakes, followed by skeleton mistakes. We do not correct both simultaneously because overly complex feedback could confuse the LLM. We first correct the linked entities since the entity information can help LLMs correct the skeleton. For example, if the user question asks for the “student with the top-ranked score”, and the database directly provides a “rank” column, the SQL can directly select data with “rank=1” without “ORDER BY score”.
4 Experiments
4.1 Settings
Datasets
We use Spider (Yu et al. 2018), Bird (Li et al. 2023b), and KaggleDBQA (Lee, Polozov, and Richardson 2021) to evaluate DAC. Spider is one of the most mainstream text-to-SQL datasets, which covers multi-table and multi-domain questions. Compared with Spider, Bird is more difficult, and databases are more complex and more near to the practical application. The difficulty of KaggleDBQA is close to that of Bird, but the scale of the database is larger [^2] We call KaggleDBQA as Kaggle for simplicity..
Metric
In this paper, following the previous works (Askari, Poelitz, and Tang 2024; Cen et al. 2024), we use execution accuracy (EX) to evaluate our method. EX refers to the proportion of data that predicted SQL execution results are the same as correct SQL execution results.
Models
We verify our methods on the instruction fine-tuning version of Llama3 (Meta 2024), Deepseek-Coder (Deep.C.) (Guo et al. 2024c), and gpt-3.5-turbo-1106 (GPT-3.5) (Ye et al. 2023). Llama3 and Deepseek-Coder are the most mainstream general and code LLMs in the current open-source LLMs, respectively. While GPT-3.5 is one of the most popular closed-source LLMs. Therefore, our experimental LLMs can well cover different application scenarios.
Implementation Details
We employ few-shot inference for each step in DAC. Following the previous work (Chang and Fosler-Lussier 2023), we use 5-shot for the inference. We use the Spider training set as the candidate demonstration pool and use BM25 to select the most similar demonstrations for each user question. For the entity linking step, we employ the labeled linking information of Spider collected by Liu et al. (2021), which is used to guide the generation format (Madaan, Hermann, and Yazdanbakhsh 2023) and can be adapted to different databases as shown in Table 1. The prompts we used can be seen in the supplementary material.
4.2 Main Experiment
Model Method Spider Bird Kaggle
Llama3-8b few-shot $70.7$ $30.4$ $24.6$
   + DAC $75.0$ $37.3$ $29.4$
Llama3-70b few-shot $80.7$ $47.0$ $31.4$
   + DAC $81.3$ $48.4$ $35.1$
Deep.C.-6.7b few-shot $74.1$ $42.4$ $21.7$
   + DAC $77.1$ $44.7$ $25.4$
Deep.C.-33b few-shot $77.0$ $48.3$ $27.0$
   + DAC $80.1$ $52.7$ $29.2$
GPT-3.5 few-shot $76.9$ $40.0$ $24.3$
   + DAC $80.6$ $44.9$ $30.9$
Table 1: The main experiment results of DAC. The baseline few-shot method uses the initial SQL as the result. The improved results led by DAC are marked in bold.
Model Method Spider Bird
gpt-3.5-turbo Self-Debug (Chen et al. 2024) $71.1\rightarrow 72.2$-
SQL-CRAFT (Xia et al. 2024) $78.2\rightarrow 79.3$-
DAC $76.9\rightarrow\bm{80.6}$ $40.0\rightarrow 44.9$
gpt-4 Self-Debug (Chen et al. 2024) $73.2\rightarrow 73.6$-
EPI-SQL (Liu and Tan 2024) $79.7\rightarrow 85.1$-
SQL-CRAFT (Xia et al. 2024) $79.7\rightarrow 85.4$ $53.3\rightarrow 55.2$
MAGIC (Askari, Poelitz, and Tang 2024) $78.6\rightarrow\bm{85.7}$-
Table 2: The comparison of DAC with other text-to-SQL correction methods on GPT-3.5 and gpt-4. Since the performance is affected by the API version, we report the performance change after using each method. We do not report the performance on KaggleDBQA since other text-to-SQL correction methods do not report the performance in their paper on this dataset. We do not adapt DAC to gpt-4 due to the high cost of the API.
The main experiment results of our method are shown in Table 1. From the table, it can be found that DAC brings performance improvement on all datasets and all models, with an average improvement of $3.7\%$, proving the effectiveness and generalization of our method.
We also compare our method with other text-to-SQL correction methods, as shown in Table 2. It can be observed that on GPT-3.5, our method achieves better performance improvement despite having a lower baseline performance, demonstrating that our method outperforms existing text-to-SQL correction methods on GPT-3.5.
Besides, from Table 1, we can also see that:
Dataset
DAC improves performance across different datasets. Compared to Spider, our method achieves more performance gains on Bird and KaggleDBQA in most settings. This is because, with more complex questions and databases, LLMs are more likely to make entity-linking mistakes or generate incorrect skeletons. Consequently, our method enhances model performance more effectively in Bird and KaggleDBQA by correcting skeleton or entity mistakes.
Model
Our method also brings significant performance improvements across different models. Compared to code-specific LLMs, the enhancement is more pronounced in general LLMs. This is because the code LLMs are fine-tuned on the code data, making its instruction following ability on non-code tasks weaker than that of the general LLMs, leading to poorer performance in entity linking and skeleton parsing.
Scale
DAC improves performance across models of different scales. The performance improvement is more significant for small-scale models compared to large-scale models. This is because small-scale models are more prone to mistakes in entities and skeletons due to their performance limitations. Our method, by correcting from the perspective of sub-tasks, effectively guides the model to generate correct results.
4.3 Ablation Study
Model Scale Method Spider Bird Kaggle
Llama3 8b DAC $75.0$ $37.3$ $29.4$
   - Entity $74.9(-0.1)$ $36.4(-0.9)$ $29.2(-0.2)$
   - Skeleton $73.3(-1.7)$ $35.3(-2.0)$ $29.0(-0.4)$
70b DAC $81.3$ $48.4$ $35.1$
   - Entity $81.2(-0.1)$ $47.1(-1.3)$ $34.9(-0.2)$
   - Skeleton $80.9(-0.4)$ $47.1(-1.3)$ $33.8(-1.3)$
Deep.C. 6.7b DAC $77.1$ $44.7$ $25.4$
   - Entity $77.0(-0.1)$ $44.1(-0.6)$ $25.4(-0.0)$
   - Skeleton $76.5(-0.6)$ $44.6(-0.1)$ $25.2(-0.2)$
33b DAC $80.1$ $52.7$ $29.2$
   - Entity $79.9(-0.2)$ $52.1(-0.6)$ $28.3(-0.9)$
   - Skeleton $79.4(-0.7)$ $51.6(-1.1)$ $28.9(-0.3)$
Table 3: The ablation experiments of DAC on: (i) Entity: remove the linked entities during the correction; (ii) Skeleton: remove the parsed skeleton during the correction.
To validate the effectiveness of each step that our method designed, we conduct ablation experiments on entity linking and skeleton parsing. The results are shown in Table 3, from which we can see that: (i) Removing either entity linking or skeleton parsing results in a performance drop, confirming the effectiveness of each step of our design; (ii) The performance drop is more significant when skeleton parsing is removed compared to entity linking, indicating that the primary performance bottleneck of the current model lies in skeleton parsing; (iii) The performance degradation is more pronounced in models with smaller scales after ablation, suggesting that errors in the results generated by smaller models are more significant, and thus DAC can bring greater performance improvements.
4.4 Analysis
Is SQL with the correct entity and skeleton the correct SQL?
Model Scale Spider Bird Kaggle
Llama3 8b $99.7$ $93.8$ $94.0$
70b $99.7$ $93.3$ $96.8$
Deep.C. 6.7b $99.7$ $96.1$ $98.9$
33b $99.7$ $94.2$ $96.0$
Table 4: The proportion of the correct SQL in the results that the linked entities and the skeleton are correct.
In the Discussion section, we propose that automated correction from the perspective of sub-tasks could not yield correct SQL results. To demonstrate the effectiveness of DAC, we calculate the proportion of correct SQL among those with correct entities and skeletons, as shown in Table 4. From the table, we can observe that: (i) The results under most settings are close to $100\%$, proving a strong consistency between automated correction from the perspectives of entities and skeletons and the correct SQL, thereby validating the effectiveness of our method; (ii) Compared to Spider, the accuracy of results on Bird and KaggleDBQA is lower, indicating that these two datasets are more challenging and require abilities beyond entity linking and skeleton parsing to produce correct SQL.
Can LLMs effectively generate entity linking information?
Method Table Column
P R F P R F
N-Gram 78.2 69.6 73.6 61.4 69.1 65.1
EtA+BERT 81.1 85.3 83.1 86.1 79.3 $82.5$
Llama3-8b 80.9 86.4 81.6 76.6 78.3 75.5
Llama3-70b $88.9$ $89.1$ $87.4$ 78.9 $83.1$ 79.0
Deep.C.-6.7b 69.4 74.6 70.0 72.8 72.6 70.3
Deep.C.-33b 79.1 82.5 79.1 $86.2$ 78.6 79.9
Table 5: The entity linking performance of different models on Spider. P, R, and F represent the macro average of accuracy, recognition rate, and F1 respectively. The results of N-Gram and EtA+BERT are reported by Liu et al. (2021) The best performances are marked in bold.
To validate the performance of LLMs in entity linking, we conduct experiments on the schema linking of the Spider dev set. The experimental results are shown in Table 5, from which we can see that: (i) In most metrics, the performance of entity linking based on LLMs is superior to that based on grams or fine-tuning methods (Liu et al. 2021), demonstrating the effectiveness of LLMs in entity linking; (ii) Compared to code-specific LLMs, general LLMs exhibit better entity-linking performance in most metrics, indicating that general-purpose LLMs are more suitable for the entity-linking task due to the better instruction following.
Is the SQL skeleton generated not based on databases better than based on databases?
Model Scale Method Spider Bird Kaggle
Llama3 8b Initial $32.0$ $8.1$ $18.2$
Parsed $\bm{35.1}$ $\bm{10.1}$ $\bm{19.5}$
70b Initial $36.1$ $11.7$ $24.3$
Parsed $\bm{37.4}$ $\bm{12.0}$ $\bm{25.2}$
Deep.C. 6.7b Initial $36.0$ $13.2$ $15.4$
Parsed $\bm{37.7}$ $\bm{18.3}$ $\bm{19.9}$
33b Initial $35.0$ $11.5$ $17.5$
Parsed $\bm{36.5}$ $\bm{16.2}$ $\bm{19.3}$
Table 6: The skeleton accuracy of initial SQL and parsed skeletons. The best performance of each model and each dataset is marked in bold.
To verify whether the performance of skeletons generated by LLMs is higher than directly generating SQL, we evaluate the accuracy of the skeletons generated on various datasets and models. The experimental results are shown in Table 6, from which we can observe: (i) Across all datasets and models, the accuracy of the parsed skeletons is higher than that of directly generated SQL, demonstrating the effectiveness of using parsed skeletons for correction; (ii) The improvement in skeleton accuracy is more significant in LLMs with smaller scales since small-scale models are less robust and more affected by disturbances in database information, leading to worse initial SQL; (iii) The accuracy of generating skeletons is lower than that of generating SQL in Table 1 because one question could correspond to multiple SQL skeletons, so the results in Table 6 can only reflect the correlation between the skeleton and the question to a certain extent.
What is the performance of DAC using oracle entities and skeletons?
Model Scale Method Spider Bird Kaggle
Llama3 8b DAC $75.0$ $37.3$ $29.4$
   + Oracle Entity $75.9(+0.9)$ $37.7(+0.4)$ $29.8(+0.4)$
   + Oracle Skeleton $75.6(+0.6)$ $37.9(+0.6)$ $30.5(+1.1)$
   + Oracle Both $76.4(+1.4)$ $38.0(+0.7)$ $30.7(+1.3)$
70b DAC $81.3$ $48.4$ $35.1$
   + Oracle Entity $81.8(+0.5)$ $48.6(+0.2)$ $37.3(+2.2)$
   + Oracle Skeleton $83.6(+2.3)$ $48.9(+0.5)$ $36.9(+1.8)$
   + Oracle Both $83.8(+2.5)$ $49.5(+1.1)$ $38.6(+3.5)$
Deepssek-Coder 6.7b DAC $77.1$ $44.7$ $25.4$
   + Oracle Entity $77.9(+0.8)$ $44.7(+0.0)$ $26.3(+0.9)$
   + Oracle Skeleton $78.9(+1.8)$ $44.8(+0.1)$ $26.5(+1.1)$
   + Oracle Both $78.9(+1.8)$ $45.4(+0.7)$ $27.2(+1.8)$
33b DAC $80.1$ $52.7$ $29.2$
   + Oracle Entity $80.7(+0.6)$ $52.7(+0.0)$ $31.1(+1.9)$
   + Oracle Skeleton $81.0(+0.9)$ $52.8(+0.1)$ $30.0(+0.8)$
   + Oracle Both $81.6(+1.5)$ $53.1(+0.4)$ $31.1(+1.9)$
Table 7: The performance of DAC with oracle entities and skeletons. Entity and Skeleton denote using the oracle entities and the oracle skeleton respectively, and Both denote using the oracle of them both.
To explore the bottlenecks in improving the performance of DAC, we conduct experiments using oracle entities and skeletons. The experimental results are shown in Table 7, from which we can observe: (i) Introducing oracle entities or skeletons further improves the performance of DAC, demonstrating that enhancing entity linking or skeleton parsing can further enhance the performance; (ii) The performance improvement using oracle skeletons is greater than using oracle entities, indicating that the performance limitation of skeleton parsing is more severe than that of entity linking; (iii) The performance improvement using the oracle entities or oracle skeletons is not significant, suggesting that direct correction by LLMs is not effective, where future work should explore more effective methods to help LLMs align the SQL to the linked entities or generated skeletons.
What is the performance of DAC on SQLs with different hardness?
Model Method Easy Medium Hard Extra
Llama3-8b baseline $87.1$ $78.0$ $59.8$ $38.0$
+ DAC $89.9$ $81.8$ $61.5$ $48.8$
Llama3-70b baseline $92.3$ $87.0$ $68.4$ $59.0$
+ DAC $94.0$ $86.5$ $69.0$ $61.4$
Deep.C.-6.7b baseline $86.7$ $83.0$ $61.5$ $44.6$
+ DAC $90.3$ $84.8$ $66.7$ $47.6$
Deep.C.-33b baseline $92.3$ $85.7$ $59.8$ $48.8$
+ DAC $95.2$ $86.8$ $62.1$ $58.4$
Table 8: Performance on questions with different hardness of the Spider dev set with and without DAC. The best performance of different settings is marked in bold.
To evaluate the performance of DAC on questions of varying difficulty, we analyze the results of the Spider dev set, which is shown in Table 8. From the table, we can observe: (i) Our method significantly improves performance across questions of different hardness, demonstrating its effectiveness on both easy and hard questions; (ii) Compared to easy questions, the performance improvement is more pronounced on harder questions, since LLMs are more likely to make mistakes in entity linking or skeleton parsing.
4.5 Case Study
Figure 4: The case study with and without DAC of the Spider dev set using Llama3-70b. The mistake in the SQL is annotated with bold.
Despite the effectiveness of DAC is demonstrated in Table 1, how our method specifically enhances text-to-SQL performance remains to be studied. Therefore, we conduct a case study, as shown in Figure 4. From the figure, we can see that, without our method, the model output misses using entities and generates an incorrect skeleton. With our method, the model generates the correct skeleton and utilizes all relevant entities, thereby enhancing text-to-SQL performance.
5 Related Works
5.1 Text-to-SQL
Research on the text-to-SQL task investigates how to generate corresponding SQL results based on given user questions and databases (Deng, Chen, and Zhang 2022). The past research is mainly based on small-scale models (Guo et al. 2019; Bogin, Gardner, and Berant 2019; Wang et al. 2020; Scholak, Schucher, and Bahdanau 2021; Dou et al. 2022) like BERT (Devlin et al. 2019) and BART (Lewis et al. 2020). Currently, the methods based on LLMs have become the mainstream of the text-to-SQL task since their brilliant performance without fine-tuning (Hong et al. 2024b). One type of approach is to fine-tune LLMs to help them better adapt to the text-to-SQL task by directly using the text-to-SQL data (Zhang et al. 2024b; Li et al. 2024b; Pourreza and Rafiei 2024), or merging the data of the tasks related to the text-to-SQL task (e.g., entity linking, knowledge generation) (Sun et al. 2024; Hong et al. 2024a). Another type of approach uses LLMs by few-shot inference, achieving significantly better performance without the overhead of fine-tuning, like in-context learning (Pourreza and Rafiei 2023; Gao et al. 2024; Chang and Fosler-Lussier 2023), self-consistency (Ni et al. 2023; Zhong et al. 2023; Lee et al. 2024), and multi-agent (Wang et al. 2024a; Xie et al. 2024).
However, it is hard for the existing methods to effectively judge the correctness of generated SQL, limiting their practical application. Therefore, we propose our method, which automatically assesses the correctness based on the decomposed sub-tasks of entity linking and skeleton parsing, thereby correcting the mistakes in the generated SQL.
5.2 Automated Correction
The automated correction approach first determines the correctness of the generated result after producing an answer, where if the answer is incorrect, it makes the necessary corrections (Pan et al. 2024). The most mainstream automated correction methods primarily focus on the code generation task, such as using feedback during training to correct errors (Ouyang et al. 2022; Feng et al. 2024; McAleese et al. 2024) or having the model judge and correct errors after inference (Chen et al. 2023, 2024; Wu et al. 2024; Zhang et al. 2024c). For the text-to-SQL task, automated correction methods pay more attention to the characteristics of the task itself, such as correcting based on execution results from the database (Li and Xie 2024; Zhong et al. 2023; Xia et al. 2024) or generating SQL-related error messages (Liu and Tan 2024; Askari, Poelitz, and Tang 2024).
However, previous work shows that LLMs struggle to directly identify SQL mistakes (Zhang et al. 2024a; Kamoi et al. 2024). Therefore, we propose DAC to provide decomposed feedback from the sub-tasks of entity linking and skeleton parsing, thereby helping correct mistakes.
6 Conclusion
In this paper, we try to alleviate the problem that LLMs perform poorly when directly correcting generated SQL for the text-to-SQL task. We first propose that correcting by decomposing the original task into sub-tasks is more effective than directly correcting the generated results. Based on this discussion, we propose DAC that decomposes the text-to-SQL task into two sub-tasks: entity linking and skeleton parsing. We ask LLMs to generate entity and skeleton information, identify inconsistencies with the initial SQL as feedback, and use this feedback for correction. Experimental results show that our method brings a $3.7\%$ performance improvement on three mainstream text-to-SQL datasets, demonstrating the effectiveness of our method.
Appendix A Prompts of DAC
The prompt of SQL Generation.
Generate a SQL to answer the question with the given schema.
Quote your answer with:
“‘sql
$<$answer sql$>$
For example:
“‘sql
{schema of demonstrations}
Question: {question of demonstrations}
“‘sql
{sql of demonstrations}
Based on the instruction and the examples, answer the following question:
“‘sql
{schema}
Question: {question}
Table 9: The prompt of SQL Generation of DAC.
The prompt of Entity Linking.
Align the tokens in the given question to the table entities or the column entities of the schema above, considering the given SQL.
Present the aligned tokens in the python format List[Dict[str, str]], where each Dict[str, str] denoting each token in the question containing the following keys:
”token”: the token in the question
”schema”: the schema entity aligned to the token
”type”: the type of the entity aligned to the token
The ”type” can be one of the following:
”tbl”: the table name
”col”: the column name
”val”: the value
”schema” and ”type” are either both null or not null at the same time.
Here are some examples.
{schema of demonstrations}
SQL: {sql of demonstrations}
Question: {question of demonstrations}
Alignments: {alignment of demonstrations}
Based on the instruction and the examples above, solve the following question:
{schema}
SQL: {sql}
Question: {question}
Alignments:
Table 10: The prompt of Entity Linking of DAC.
The prompt of Skeleton Parsing.
Hallucinate a SQL to answer the question.
Quote your answer with:
“‘sql
$<$answer sql$>$
For example:
Question: {question of demonstrations}
“‘sql
{sql of demonstrations}
Based on the instruction and the examples, answer the following question:
Question: {question}
Table 11: The prompt of Skeleton Parsing of DAC.
The prompt of Correction with the entity feedback.
“‘sql
{schema}
Fix the sql ”{sql}” to answer the question ”{question}” based on the above database and the alignment.
Present your sql in the format:
“‘sql
$<$your sql$>$
It should be noticed that {notification}. Your sql must contain the tables and columns mentioned by the question.
Table 12: The prompt of Correction of DAC with the entity feedback. The format of “{notification}” is like “<<<tables or columns>>> are mentioned by the question”.
The prompt of Correction with the skeleton feedback.
“‘sql
{schema}
Fix the sql ”{sql}” to answer the question ”{question}” with the above schema.
Present your sql in the format:
“‘sql
$<$your sql$>$
It should be noticed that the SQL skeleton could be like ”{skeleton}”, where each ’_’ can only be replaced with one single table, column or value.
Table 13: The prompt of Correction of DAC with the skeleton feedback.
The prompts of DAC is shown in Table 9, Table 10, Table 11, Table 12, and Table 13. The format of “{schema}” in our prompts is the same as the definition in (Chang and Fosler-Lussier 2023).
Appendix B Error Analysis
TableColumnSkeletonExecution00200200200200400400400400195.2195.2195.2195.2403.3403.3403.3403.3497.2497.2497.2497.2133.5133.5133.5133.5173.6173.6173.6173.6370.9370.9370.9370.9458.3458.3458.3458.342.242.242.242.2Error Type#Examplebaseline+DAC
Figure 5: The error analysis with and without DAC. #Example denotes the average error examples across both models (Llama3, Deepseek-Coder) and all three datasets (Spider, Bird and KaggleDBQA). The table and column errors denote that the generated SQL does not have corresponding tables and columns of the correct SQL. The skeleton error denotes not having the correct skeleton. The execution error denotes that the generated SQL is not able to be executed. There could be one example that has multiple errors.
To analyze the limitations of DAC and inspire future research, we compile statistics on the errors produced by our method. We categorize these errors into table errors, column errors, skeleton errors, and execution errors. The specific statistical results are shown in Figure 5, from which we can observe: (i) Our method significantly reduces errors across all categories, demonstrating its effectiveness; (ii) The primary error remains in skeleton generation due to that one user question could correspond to multiple different SQLs; (iii) Besides skeleton errors, column errors are also significant, indicating that effectively guiding LLMs to retrieve and use relevant columns remains under discovery.